OpenStudy (anonymous):

Consider the functions f(x) = 2x2 + 1 and g(x) = x + 2. What is the value of f[g(3)]? Answer 11 21 101 51

OpenStudy (anonymous):

51

OpenStudy (anonymous):

3 + 2 = g(3) = 5 f(5) = 2 * 5 * 5 + 1 = 51